### v2

# iclr2024_conference

[…]
iclr2024_conference
Appendix: Decoupling learning from negative and positive feedback in preference optimization
Antiquus S. Hippocampus, Natalia Cerebro & Amelie P. Amygdale
Department of Computer Science
Cranberry-Lemon University
Pittsburgh, PA 15213, USA
{hippo,brain,jen}@cs.cranberry-lemon.edu
&Ji Q. Ren & Yevgeny LeNet
Department of Computational Neuroscience
University of the Witwatersrand
Joburg, South Africa
{robot,net}@wits.ac.za
[…]
Affiliation
Address
email
Use footnote for providing further information about author (webpage, alternative address)—not for acknowledging funding agencies. Funding acknowledgements go at the end of the paper.
Appendix A Hyperparameter tuning
PMPO introduces two hyperparameters, $\alpha$ and $\beta$. However, we found them relatively easy to tune in practice based on the following guidelines:
  1. $\boldsymbol{\alpha}$: This parameter reflects the relative importance of positive and negative feedback. When both are available, $\alpha=0.5$ is a reasonable starting point if both types of feedback are equally reliable. Otherwise, α can be adjusted to reflect the confidence in each feedback type. In cases with only one type of feedback, α is naturally determined by the data.
  2. $\boldsymbol{\beta}$: This parameter controls the influence of the prior or reference policy. Our experiments suggest that a higher β is generally beneficial when learning from negative feedback such that the more is contribution of negative feedback to the policy update (lower $\alpha$) the more $\beta$ should be. This can also be motivate from the derivations that KL term is appear as the result of learning from negative feedback. We also find that as long as $\beta$ is large enough, the algorithm is fairly insensitive to the exact choice of the parameter.
Appendix B Computing the Kullback-Leibler divergence
learning from negative feedback is only feasible with a correct KL term and access to a reference policy $\pi_{ref}$. When logits are available using an exact KL computation is ideal. In the absence of logits (Experiment 5.3), we relied on an abundance of unlabelled data to estimate it. This enables us to learn only from negative feedback without access to positive feedback as suggested by our derivations.
Subsequently, we present the derivations for computing KL divergence for LLMs with autoregressive policies in our experiments.
Let’s consider a single two-token generation $(y_{1},y_{2})$ sampled autoregressively from a prompt $x$. We can factorize the joint distribution associated with such a generation as follows:
$\pi(y_{1},y_{2}|x)=\pi^{1}(y_{1}|x)\pi^{2}(y_{2}|x,y_{1})$ (1)
Using this factorization for both $\pi_{\theta}$ and $\pi_{\text{ref}}$ in our KL regularizer, we compute the following:
$\operatorname*{\mathrm{KL}}(\pi_{\text{ref}}\|\pi_{\theta})$ $=\mathbb{E}_{\pi_{\text{ref}}}\log\frac{\pi_{\text{ref}}}{\pi_{\theta}}$ (2)
$=\mathbb{E}_{\pi_{\text{ref}}^{1}}\mathbb{E}_{\pi_{\text{ref}}^{2}}\log\frac{\pi_{\text{ref}}^{1}\ \pi_{\text{ref}}^{2}}{\pi_{\theta}^{1}\ \pi_{\theta}^{2}}$ (3)
$=\mathbb{E}_{\pi_{\text{ref}}^{1}}\mathbb{E}_{\pi_{\text{ref}}^{2}}\log\frac{\pi_{\text{ref}}^{1}}{\pi_{\theta}^{1}}+\mathbb{E}_{\pi_{\text{ref}}^{1}}\mathbb{E}_{\pi_{\text{ref}}^{2}}\log\frac{\pi_{\text{ref}}^{2}}{\pi_{\theta}^{2}}$ (4)
$=\mathbb{E}_{\pi_{\text{ref}}^{1}}\log\frac{\pi_{\text{ref}}^{1}}{\pi_{\theta}^{1}}+\mathbb{E}_{\pi_{\text{ref}}^{1}}\mathbb{E}_{\pi_{\text{ref}}^{2}}\log\frac{\pi_{\text{ref}}^{2}}{\pi_{\theta}^{2}},$ (5)
where the first term can drop the expectation with respect to $\mathbb{E}_{\pi_{\text{ref}}^{2}}$ since the integrand $\log\frac{\pi_{\text{ref}}^{1}}{\pi_{\theta}^{1}}$ does not depend on $y_{2}$. The first term can be computed analytically because $\pi_{\text{ref}}^{1}$ and $\pi_{\theta}^{1}$ are simply categorical distributions over the entire vocabulary of tokens. The second term, however, is problematic due to the outer expectation $\mathbb{E}_{\pi_{\text{ref}}^{1}}$, which requires us to integrate a KL over all possible values of $y_{1}$. While we can easily compute this KL for any one value of $y^{1}$, the integration would be unwieldy and certainly becomes intractable as we extend this to longer sequences $y_{3}$, $y_{4}$, etc. Luckily, during training we obtain samples $\tilde{y}_{1},\tilde{y}_{2}\sim\pi_{\text{ref}}$ and so $\tilde{y}_{1}\sim\pi_{\text{ref}}^{1}$, which allows us to use a single-sample Monte Carlo unbiased estimate of the second term such that:
$\approx\mathbb{E}_{\pi_{\text{ref}}^{1}}\log\frac{\pi_{\text{ref}}^{1}}{\pi_{\theta}^{1}}+\mathbb{E}_{\pi_{\text{ref}}^{2}}\log\frac{\pi_{\text{ref}}^{2}(\cdot|x,\tilde{y}_{1})}{\pi_{\theta}^{2}(\cdot|x,\tilde{y}_{1})}.$ (6)
Extending this to sequences of length $L$, and dropping the superscripts on the policies as they are autoregressive, we obtain the following approximation:
$\operatorname*{\mathrm{KL}}(\pi_{\text{ref}}\|\pi_{\theta})$ $\approx\mathbb{E}_{\pi_{\text{ref}}}\log\frac{\pi_{\text{ref}}(\cdot|x)}{\pi_{\theta}(\cdot|x)}+\sum_{i=1}^{L-1}\mathbb{E}_{\pi_{\text{ref}}}\log\frac{\pi_{\text{ref}}(\cdot|x,\tilde{y}_{1:i})}{\pi_{\theta}(\cdot|x,\tilde{y}_{1:i})}.$ (7)
where $\tilde{y}_{1:L}\sim\pi_{\text{ref}}$.
Appendix C Fundamental results
This section explains precisely why the EM method is a sound approach to optimization. We will present the case of a discrete function. This section relies on classical results in discrete optimization and shows how one can build a simple strictly improving algorithm that maximises a discrete function. For simplicity and clarity, we will optimize a function $f\in\mathbb{R}^{\mathcal{S}}$ mapping elements of $\mathcal{S}$ to real numbers, where $\mathcal{S}$ is a finite set. More precisely, our goal is to find a distribution $\delta\in\Delta_{\mathcal{S}}$ (parameterized policy in RL) that maximises the expectation of $f$ under $\delta$ (expected value function in RL):
$\mathbb{E}_{s\sim\delta}[f(s)]=\sum_{s\in\mathcal{S}}\delta(s)f(s).$
We recall that a discrete probability distribution $\delta\in\Delta_{\mathcal{S}}$ can be identified as a positive real function $\delta\in\mathbb{R}_{+}^{\mathcal{S}}$ verifying:
$\sum_{s\in\mathcal{S}}\delta(s)=1.$
To find a good distribution to maximise $\mathbb{E}_{s\sim\delta}[f(s)]$, the algorithm relies on the following results:
  • Starting from a distribution $\eta\in\Delta_{\mathcal{S}}$, there exists a unique closed-form argmaximum $\delta^{*}$ of the regularised expectation $\mathbb{E}_{s\sim\delta}[f(s)]-\tau\text{KL}(\delta\;||\;\eta)$, where $\tau\in\mathbb{R}^{*}_{+}$ is a strictly positive real number. We have $\delta^{*}=\frac{\eta(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}$.
  • Unless $\delta^{*}=\eta$, we have a strict improvement between $\delta^{*}$ and $\eta$:
    $\mathbb{E}_{s\sim\delta^{*}}[f(s)]>\mathbb{E}_{s\sim\eta}[f(s)].$
We prove these results in the main paper. Those results implies that the following algorithm that starts at $\delta_{0}=\eta$ and computes:
$\forall k\in\mathbb{N},\delta_{k+1}=\frac{\delta_{k}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}\delta_{k}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))},$
is strictly monotonically improving until there is $K\in\mathbb{N}$ such that $\delta_{K+1}=\delta_{K}$:
$\mathbb{E}_{s\sim\delta_{0}}[f(s)]<\mathbb{E}_{s\sim\delta_{1}}[f(s)]<\dots<\mathbb{E}_{s\sim\delta_{K}}[f(s)]=\mathbb{E}_{s\sim\delta_{K+1}}[f(s)].$
In practice, we have a set of learnable weights $\theta$ to parameterize a distribution $q_{\theta}$ in order to fit $\delta_{k+1}$ and another set of fixed weights $\mu$ to parameterize a distribution $q_{\mu}=\delta_{k}$. Then, to fit $\delta_{k+1}$ the idea is to minimize the following KL divergence (this is the maximisation step):
$\mathcal{L}(\theta)$ $=\text{KL}(\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}\;||\;q_{\theta}),$
$=\mathbb{E}_{s\sim\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}\left[\log\left(\frac{\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}{q_{\theta}}\right)\right],$
$=-\mathbb{E}_{s\sim\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}\left[\log\left(q_{\theta}\right)\right]+\mathbb{E}_{s\sim\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}\left[\log\left(\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}\right)\right]$
The term $\mathbb{E}_{s\sim\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}\left[\log\left(\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}\right)\right]$ does not depend on $\theta$ so is irrelevant in the minimization. Using the re-weighting formula $\mathbb{E}_{s\sim\delta}[f(s)]=\mathbb{E}_{s\sim\eta}[\frac{\delta(s)}{\eta(s)}f(s)]$, the minimization problem is equivalent to:
$\mathcal{L}(\theta)$ $=-\mathbb{E}_{s\sim\frac{q_{\mu}(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}\left[\log\left(q_{\theta}\right)\right],$
$=-\mathbb{E}_{s\sim q_{\mu}}\left[\frac{\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}\log\left(q_{\theta}\right)\right],$
As $\sum_{s^{\prime}\in\mathcal{S}}q_{\mu}(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))$ is a constant, this is equivalent to minimizing:
$\mathcal{L}(\theta)=-\mathbb{E}_{s\sim q_{\mu}}[\exp(\tau^{-1}f(s))\log\left(q_{\theta}(s)\right)].$
C.1 Existence and uniqueness of the regularized argmaximum
For completeness, we briefly recall the proof of existence and uniqueness of the argmaximum of the following regularized criterion that can also be found in the work of rafailov2023direct:
$\mathcal{L}_{\tau}(\delta)$ $=\mathbb{E}_{s\sim\delta}[f(s)]-\tau\text{KL}(\delta\;||\;\eta),$
$=\sum_{s\in\mathcal{S}}\delta(s)f(s)-\tau\text{KL}(\delta\;||\;\eta).$
Now, if we define the softmax probability $\delta^{*}\in\Delta_{\mathcal{S}}$ as:
$\forall s\in\mathcal{S},\delta^{*}(s)=\frac{\eta(s)\exp(\tau^{-1}f(s))}{\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))},$
then, under the previous definitions, we have the following results:
$\delta^{*}=\operatorname*{\mathrm{arg~{}max}}_{\delta\in\Delta_{\mathcal{S}}}\mathcal{L}_{\tau}(\delta)$
Proof.
$\frac{\mathcal{L}_{\tau}(\delta)}{\tau}$ $=\sum_{s\in\mathcal{S}}\delta(s)\frac{f(s)}{\tau}-\text{KL}(\delta\;||\;\eta),$
$=\sum_{s\in\mathcal{S}}\delta(s)\frac{f(s)}{\tau}-\sum_{s\in\mathcal{S}}\delta(s)\log\big{(}\frac{\delta(s)}{\eta(s)}\big{)},$
$=\sum_{s\in\mathcal{S}}\delta(s)\big{(}\frac{f(s)}{\tau}-\log\big{(}\frac{\delta(s)}{\eta(s)}\big{)}\big{)},$
$=\sum_{s\in\mathcal{S}}\delta(s)\big{(}\log\big{(}\exp(\tau^{-1}f(s))\big{)}-\log\big{(}\frac{\delta(s)}{\eta(s)}\big{)}\big{)},$
$=\sum_{s\in\mathcal{S}}\delta(s)\big{(}\log\big{(}\frac{\eta(s)\exp(\tau^{-1}f(s))}{\delta(s)}\big{)}\big{)},$
$=\sum_{s\in\mathcal{S}}\delta(s)\big{(}\log\big{(}\frac{\eta(s)\exp(\tau^{-1}f(s))\frac{\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}{\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}{\delta(s)}\big{)}\big{)},$
$=\sum_{s\in\mathcal{S}}\delta(s)\big{(}\log\big{(}\frac{\frac{\eta(s)\exp(\tau^{-1}f(s))}{\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}}{\delta(s)}\big{)}\big{)}+\sum_{s\in\mathcal{S}}\delta(s)\log\big{(}\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))\big{)},$
$=\sum_{s\in\mathcal{S}}\delta(s)\big{(}\log\big{(}\frac{\delta^{*}(s)}{\delta(s)}\big{)}\big{)}+\log\big{(}\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))\big{)},$
$=-\text{KL}(\delta\;||\;\delta^{*})+\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}.$
By definition of the KL, we now that $\delta^{*}=\operatorname*{\mathrm{arg~{}max}}_{\delta\in\Delta_{\mathcal{S}}}\bigg{[}-\text{KL}(\delta\;||\;\delta^{*})\bigg{]}$ and as:
$-\text{KL}(\delta\;||\;\delta^{*})=\frac{\mathcal{L}_{\tau}(\delta)}{\tau}-\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}$ (8)
where $\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}$ is a constant (does not depend on $\delta$) and $\tau$ a positive multiplicative term, then $-\text{KL}(\delta\;||\;\delta^{*})$ and $\mathcal{L}_{\tau}(\delta)$ share the same argmaximum. This concludes the proof. ∎
The fact that we have:
$\delta^{*}=\frac{\eta(\cdot)\exp(\tau^{-1}f(\cdot))}{\sum_{s^{\prime}\in\mathcal{S}}\eta(s^{\prime})\exp(\tau^{-1}f(s^{\prime}))}=\operatorname*{\mathrm{arg~{}max}}_{\delta\in\Delta_{\mathcal{S}}}\mathcal{L}_{\tau}(\delta),$
implies by simply replacing $\delta$ by $\delta^{*}$ in equation (8) that:
$-\text{KL}(\delta^{*}\;||\;\delta^{*})=\frac{\mathcal{L}_{\tau}(\delta^{*})}{\tau}-\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}.$
As $\text{KL}(\delta^{*}\;||\;\delta^{*})=0$, we have:
$\mathcal{L}_{\tau}(\delta^{*})=\max_{\delta\in\Delta_{\mathcal{S}}}\mathcal{L}_{\tau}(\delta)=\tau\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}.$ (9)
This result is often expressed in term of an inequality that says that the logsumexp is a majorant of the regularized expectation:
$\forall\delta\in\Delta_{\mathcal{S}},\tau\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}\geq\sum_{s\in\mathcal{S}}\delta(s)f(s)-\tau\text{KL}(\delta\;||\;\eta).$ (10)
C.2 Proof of Improvement.
We use the same notations as in the previous section, and our goal is to show that using the distribution $\delta^{*}$ instead of $\eta$ strictly increases the expected value of our function $f$:
$\mathbb{E}_{s\sim\delta^{*}}[f(s)]>\mathbb{E}_{s\sim\eta}[f(s)],$
unless $\delta^{*}=\eta$. This means that we can confidently replace $\eta$ by $\delta^{*}$ for the next iteration of the algorithm.
Proof.
From Eq.(9), we have:
$\tau\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}=\mathcal{L}_{\tau}(\delta^{*})=\mathbb{E}_{s\sim\delta^{*}}[f(s)]-\tau\text{KL}(\delta^{*}\;||\;\eta).$
Using Jensen inequality we have:
$\tau\log\big{(}\sum_{s\in\mathcal{S}}\eta(s)\exp(\tau^{-1}f(s))\big{)}$ $\geq\tau\sum_{s\in\mathcal{S}}\eta(s)\log\big{(}\exp(\tau^{-1}f(s))\big{)},$
$=\tau\sum_{s\in\mathcal{S}}\eta(s)\tau^{-1}f(s),$
$=\sum_{s\in\mathcal{S}}\eta(s)f(s)=\mathbb{E}_{s\sim\eta}[f(s)].$
This implies that:
$\mathbb{E}_{s\sim\delta^{*}}[f(s)]-\tau\text{KL}(\delta^{*}\;||\;\eta)$ $\geq\mathbb{E}_{s\sim\eta}[f(s)].$
As $\tau\text{KL}(\delta^{*}\;||\;\eta)$ is strictly positive unless $\delta^{*}=\eta$, we conclude that we have strict improvement of the algorithm unless $\delta^{*}=\eta$ which means in this case that the method has converged. ∎
Appendix D Additional Large Language Model Experiments
The pre-trained reward model used for our LLM experiments is not perfect and is subject to reward hacking (skalse2022defining). Generally, reward models only capture superficial features of the generation such as the length or the formatting. Therefore, instead of only looking at the reward scores, we need other ways to evaluate the quality of our policy, that are independent from the reward bias. We choose to use the LMSYS arena hard benchmark (li2024crowdsourced) as an evaluation set. It is composed of $500$ prompts. We decided to evaluate 3 different runs, PMPO-R (only reject), PMPO-AR (accept and reject) and PMPO-A (only accept) against the shared initial checkpoint. Each of the data-point in Fig. 1 corresponds to the win-rate ratio of a method at a given training step against the initial checkpoint. The win-rate ratio between two checkpoints is obtained by scraping the $500$ prompts of the evaluation set to obtain $500$ generations for each checkpoint. Then each prompt and couple of generations is sent to an LLM judge (gpt4 in this case) which decides what is the best generation for this given prompt. The win-ratio is simply the number of times one checkpoint wins over the other. In Fig. 1, we can see that PMPO-AR is the quickest to "hack the reward", it reaches a good performance but then in the middle of training its start hacking the reward and learns a pathological behaviour that makes it performs worse on the independent benchmark. This phenomenon has been observed consistently in RLHF.
Figure 1:
Appendix E Link Between IPO and PMPO
In this section, we draw a parallel between the IPO and the PMPO losses. For a dataset of triplets $(x^{i},y^{i}_{a},y^{i}_{r})_{i=1}^{N}$, the IPO loss is:
$\mathcal{L}_{\texttt{IPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[-\log\left(\pi_{\theta}(y^{i}_{a}|x^{i})\right)+\log\left(\pi_{\theta}(y^{i}_{r}|x^{i})\right)+\beta\left(\log\left(\frac{\pi_{\theta}(y^{i}_{a}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{a}|x^{i})}\right)-\log\left(\frac{\pi_{\theta}(y^{i}_{r}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{r}|x^{i})}\right)\right)^{2}\right].$
The IPO loss is composed of two terms the policy optimisation term:
$\mathcal{P}_{\texttt{IPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}-\log\left(\pi_{\theta}(y^{i}_{a}|x^{i})\right)+\log\left(\pi_{\theta}(y^{i}_{r}|x^{i})\right),$
and the policy regularisation term:
$\mathcal{R}_{\texttt{IPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[\left(\log\left(\frac{\pi_{\theta}(y^{i}_{a}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{a}|x^{i})}\right)-\log\left(\frac{\pi_{\theta}(y^{i}_{r}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{r}|x^{i})}\right)\right)^{2}\right].$
When $\alpha=\frac{1}{2}$, the PMPO loss can be written as:
$\mathcal{L}_{\texttt{PMPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[-\frac{1}{2}\log\left(\pi_{\theta}(y^{i}_{a}|x^{i})\right)+\frac{1}{2}\log\left(\pi_{\theta}(y^{i}_{r}|x^{i})\right)+\beta\texttt{KL}(\pi_{\texttt{ref}}(.|x^{i})||\pi_{\theta}(.|x^{i}))\right].$
The PMPO loss is also composed of two terms the policy optimisation term:
$\mathcal{P}_{\texttt{PMPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}-\frac{1}{2}\log\left(\pi_{\theta}(y^{i}_{a}|x^{i})\right)+\frac{1}{2}\log\left(\pi_{\theta}(y^{i}_{r}|x^{i})\right),$
and the policy regularisation term:
$\mathcal{R}_{\texttt{PMPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[\texttt{KL}(\pi_{\texttt{ref}}(.|x^{i})||\pi_{\theta}(.|x^{i}))\right].$
Therefore, the policy optimisation terms of the IPO $\mathcal{P}_{\texttt{IPO}}(\theta)$ and PMPO $\mathcal{P}_{\texttt{PMPO}}(\theta)$ losses are identical at a constant factor. Now we are going to create a connection between the IPO and PMPO regularisation terms when $y^{i}_{a}$ and $y^{i}_{r}$ are sampled from $\pi_{\texttt{ref}}(.|x^{i})$. The first thing to remark is that $\left(\log\left(\frac{\pi_{\theta}(y^{i}_{a}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{a}|x^{i})}\right)-\log\left(\frac{\pi_{\theta}(y^{i}_{r}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{r}|x^{i})}\right)\right)^{2}$ is an unbiased estimate of $E_{Y,Y^{\prime}\sim\pi_{\texttt{ref}}(.|x^{i})}\left[\left(\log\left(\frac{\pi_{\theta}(Y|x^{i})}{\pi_{\texttt{ref}}(Y|x^{i})}\right)-\log\left(\frac{\pi_{\theta}(Y^{\prime}|x^{i})}{\pi_{\texttt{ref}}(Y^{\prime}|x^{i})}\right)\right)^{2}\right]$. Then, we remind the reader that the variance of a random variable $X$ under distribution $\mu$ verifies:
$\texttt{VAR}_{X\sim\mu}[X]=\frac{1}{2}\mathbb{E}_{X,X^{\prime}\sim\mu}[(X-X^{\prime})^{2}],$
where $X$ and $X^{\prime}$ are independent variables with distribution $\mu$. This means that $\left(\log\left(\frac{\pi_{\theta}(y^{i}_{a}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{a}|x^{i})}\right)-\log\left(\frac{\pi_{\theta}(y^{i}_{r}|x^{i})}{\pi_{\texttt{ref}}(y^{i}_{r}|x^{i})}\right)\right)^{2}$ is an unbiased estimate of $2\texttt{VAR}_{Y\sim\pi_{\texttt{ref}}(.|x^{i})}\left[\log\left(\frac{\pi_{\theta}(Y|x^{i})}{\pi_{\texttt{ref}}(Y|x^{i})}\right)\right]$.
Therefore the expectation (over $\pi_{\texttt{ref}}$) of the IPO regularization term is:
$\mathbb{E}_{\pi_{\texttt{ref}}}\left[\mathcal{R}_{\texttt{IPO}}(\theta)\right]=\frac{2}{N}\sum_{i=1}^{N}\texttt{VAR}_{Y\sim\pi_{\texttt{ref}}(.|x^{i})}\left[\log\left(\frac{\pi_{\theta}(Y|x^{i})}{\pi_{\texttt{ref}}(Y|x^{i})}\right)\right].$
As $\texttt{VAR}_{X\sim\mu}[X]=\texttt{VAR}_{X\sim\mu}[-X]$, we also have:
$\mathbb{E}_{\pi_{\texttt{ref}}}\left[\mathcal{R}_{\texttt{IPO}}(\theta)\right]=\frac{2}{N}\sum_{i=1}^{N}\texttt{VAR}_{Y\sim\pi_{\texttt{ref}}(.|x^{i})}\left[\log\left(\frac{\pi_{\texttt{ref}}(Y|x^{i})}{\pi_{\theta}(Y|x^{i})}\right)\right].$
This is in contrast with the regularization term of PMPO:
$\mathcal{R}_{\texttt{PMPO}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[\texttt{KL}(\pi_{\texttt{ref}}(.|x^{i})||\pi_{\theta}(.|x^{i}))\right]=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{Y\sim\pi_{\texttt{ref}}(.|x^{i})}\left[\log\left(\frac{\pi_{\texttt{ref}}(Y|x^{i})}{\pi_{\theta}(Y|x^{i})}\right)\right].$
So the difference between PMPO and IPO is that PMPO will minimize the expectation of the log ratio between the reference policy and the online policy whereas IPO will minimize the variance of the same quantity.
Appendix F Derivation from the perspective of RL as inference
Consider a decision-making scenario characterized by the following elements:
Stationary State Distribution: A stationary distribution $\mu(x)$ describes the probability of encountering different states or contexts $x$. We typically have access to samples from this distribution.
Policy: A policy $\pi(y|x)$ dictates the probability of selecting an action/outcome $y$ given a state $x$.
Preference Information: We have access to knowledge for driving the likelihood of preference or dis-preference toward actions taken by a policy $\pi(y|x)$ according to a known or latent objective. This information takes one of or these forms:
  • Preference Probabilities $q(I=1|y\sim\pi,x\sim\mu)$, where $I$ is a binary indicator with $I=1$ representing preference. when only a reward or q-function is available this probability can be derived as $q(I=1|y\sim\pi,x\sim\mu)\propto\exp(Q(y,x)/\eta)$.
  • Dis-preference Probabilities $q(I=0|y\sim\pi,x\sim\mu)$, where $I=0$ denotes dis-preference. when only a reward or q-function is available this probability can be derived as $q(I=1|y\sim\pi,x\sim\mu)\propto\exp(-Q(y,x)/\eta)$.
Objective: derive an optimal policy to give a highest probability to the preferred outcomes:
[…]
Appendix G Background:Maximise Preferred Outcomes
We start with the stated objective:
[…]
This objective can be solved by EM algorithm to iterativly creating a tight lower bound on the current estimate of the objective $logp_{\texttt{ref}}(I=1)$ and then we optimise the objective. By repeating these two steps it’s guaranteed that algorithm will converge. In order to do so we use the following identity:
$\log p(X)=\int q(Z)\log\frac{p(X,Z)}{q(Z)}+\operatorname*{\mathrm{KL}}(q(Z)|p(Z|X))$
where the first term is a lower bound because $\operatorname*{\mathrm{KL}}$ is always positive. Now considering the states $x$ and outcome $y$ as latent variables:
$\log p_{\theta}(I=1)=\int\mu(x)\int q(y|x)\log\frac{p_{\theta}(I=1,y|x)}{q(y|x)}+\int\mu(x)\operatorname*{\mathrm{KL}}(q(y|x)|p_{\theta}(y|I=1,x))$
G.1 E-Step
In E-step our goal is to choose the variational distribution $q(y|x)$ such that lower bound on
$\log p_{\texttt{ref}}(I=1)$
is as tight as possible which is the case when $\operatorname*{\mathrm{KL}}$ term in equation above is zero at current estimate of policy $\pi_{\texttt{ref}}$. This simply leads to $q(y|x)=p_{\texttt{ref}}(y|I=1,x)$ or
$q(y|x)=\frac{\pi_{\texttt{ref}}(y|x)p_{\texttt{ref}}(I=1|x,y)}{p_{\texttt{ref}}(I=1|x)}$
where $p_{\texttt{ref}}(I=1|x)=\int\pi_{\texttt{ref}}(y|x)p_{\texttt{ref}}(I=1|x,y)$ is a normalizer for a given state $x$. Now in some cases such as data collected from human feedback we directly have access to sampled $q(y|x)$ according to equation above where if a $y_{a}$ is accepted $q(y|x)$ would be a distribution that for a state $x$ gives probability of 1 to $y_{a}$ and zero to every other samples, i.e, $q(y_{a}\sim\pi_{\texttt{ref}}|x)=1$ and $q(y_{!a}\sim\pi_{\texttt{ref}}|x)=0$. In case we have multiple ranked outcomes for a $x$ we again can simply assign probabilities accordingly and obtain $q(y|x)$ immediately.
Relation to Reinforcement Learning
E-step of preference objective coincides with KL regularized RL objective when a reward function or Q-function is given instead of preferences. In such case we can estimate preference function $p_{\texttt{ref}}(I=1,x,y)$ by using a transformation on reward function. Now if we choose $p_{\texttt{ref}}(I=1,x,y)\propto\exp(\frac{Q(x,y)}{\eta})$ and substituting this in expected KL term above we obtain the $\operatorname*{\mathrm{KL}}$ regularised RL objective, i.e,
$\max_{q}\int\mu(x)\int q(y|x)Q(y,x)-\eta\int\mu(x)\operatorname*{\mathrm{KL}}(q(y|x)|\pi_{\texttt{ref}}(y|x))$
This motivates the original preference optimization objective and implies that this objective can be seen as optimising for a latent Q-function or reward.
G.2 M-Step
In E-step we found non-parametric variational distributions $q(y|x)$ for $x\sim\mu(x)$ that gives higher probability to preferred actions sampled from $\pi_{\texttt{ref}}$. E step can be seen as sample based policy improvement $\pi_{\texttt{ref}}$ wrt preferences. In M-step we optimise the lower bound to obtain a new distribution, i.e,
$\max_{\theta}\int\mu(x)\int q(y|x)\log\frac{p_{\theta}(I=1,y|x)}{q(y|x)}$
where
$q(y|x)=\frac{\pi_{\texttt{ref}}(y|x)p(I=1|x,y)}{\int\pi_{\texttt{ref}}(y|x)p(I=1|y,x)}$
according to derivations in E-step. after rearranging the terms and removing the terms that do not depends on $\theta$ we get
$\min_{\theta}\int\mu(x)\operatorname*{\mathrm{KL}}(q(y|x)|\pi_{\theta}(y|x))$
which effectively is a weighted maximum likelihood objective to fit improved non-parametric policies. In practice we also add a constraint to avoid moving too far from from the $\pi_{\texttt{ref}}$, i.e,
$\min_{\theta}\int\mu(x)\operatorname*{\mathrm{KL}}(q(y|x)|\pi_{\theta}(y|x))+\beta\int\mu(x)\operatorname*{\mathrm{KL}}(\pi_{\texttt{ref}}(y|x)|\pi_{\theta}|(y|x))$
Note that this constraint can be motivated from different angles such as a prior on $\theta$ or partial EM and multi objective perspective. In terms of the direction of KL, this direction of KL is aligned with the direction of KL for fitting the distributions.
Appendix H Incorporating dis-preferred outcomes for policy optimization
In the previous section we showed how optimising for preferred outcomes can lead us to useful policy update rules. In order to motivate this objective further, we showed that typical reinforcement learning objective coincides with the objective in E-Step. However we only incorporated preferred outcomes $q(I=1)$ to update the policy. Now we ask this question how we can incorporate dis-preferred outcomes to directly optimize the policy? In order to do that we first write down the policy optimisation objective we derived in terms of preferences:
$\min_{\theta}\int\mu(x)\operatorname*{\mathrm{KL}}(q(y|x)|\pi_{\theta}(y|x))$
where
$q(y|x)=\frac{\pi_{\texttt{ref}}(y|x)p(I=1|x,y)}{\int\pi_{\texttt{ref}}(y|x)p(I=1|y,x)}.$
We also know that $p(I=1|x,y)=1-p(I=0|x,y)$. This is correct as $p(I)$ is a probability function over a binary random variable $I$. Therefore the variational distribution can be re-written in terms of dis preferences, i.e,
$q(y|x)=\frac{\pi_{\texttt{ref}}(y|x)(1-p(I=0|x,y))}{\int\pi_{\texttt{ref}}(y|x)(1-p(I=0|x,y))}.$
substituting this in the KL minimization term above we get:
$\min_{\theta}-\int\mu(x)\int\frac{1}{Z(x)}\pi_{\texttt{ref}}(y|x)\log\pi_{\theta}(y|x)+\int\mu(x)\int\frac{1}{Z(x)}\pi_{\texttt{ref}}(y|x)p(I=0|x,y)\log\pi_{\theta}(y|x)$
where $Z(x)=\int\pi_{\texttt{ref}}(y|x)(1-p(I=0|x,y))$ is a normalizer term. ‘Now Now we can write the final policy update rule in terms of dis-preferences:
$\min_{\theta}-\int\mu(x)\operatorname*{\mathrm{KL}}(t(y|x)|\pi_{\theta}(y|x))+\beta\int\mu(x)\operatorname*{\mathrm{KL}}(\pi_{\texttt{ref}}(x|y)|\pi_{\theta}(x|y))$
where $t(y|x)\propto\pi_{\texttt{ref}}(y|x)p(I=0|x,y)$ and $\beta$ is a positive tuning parameters that . $\beta$ must be sufficiently large otherwise as we will show in the experiment, it wont work in practice. This update rule intuitively goes away from dis-preferences while stay close to the reference policy.
Appendix I Final update rule: leveraging both preferences and dis-preferences
After putting things together we get the following update rule that maximises the preferred outcomes and minimizes the preferred outcomes, i.e,
$\max_{\theta}\alpha\log q_{\theta}(I=1)+(1-\alpha)\log(1-q_{\theta}(I=0))$
After repeating the EM steps to optimise this objective we get the following update rule:
$\pi_{\operatorname*{\mathrm{imp}}}=\operatorname*{\mathrm{arg~{}min}}_{\theta}\alpha\int\mu(x)\operatorname*{\mathrm{KL}}(q(y|x)|\pi_{\theta}(y|x))+\beta\int\mu(x)\operatorname*{\mathrm{KL}}(\pi_{\operatorname*{\mathrm{ref}}}|\pi_{\theta})-(1-\alpha)\int\mu(x)\operatorname*{\mathrm{KL}}(t(y|x)|\pi_{\theta}(y|x))$
where q is a distribution modified wrt preference probabilities resulting from $\pi_{\texttt{ref}}$ weighted by preference probabilities, i.e,
$q(y|x)\propto\pi_{\texttt{ref}}(y|x)p(I=1|x,y)$
or in case of access to q-function or reward function
$q(y|x)\propto\pi_{\texttt{ref}}(y|x)\exp(\frac{Q(x,y)}{\eta})$
and t is a distribution modified wrt dis-preference probabilities resulting from $\pi_{\texttt{ref}}$ weighted by dis-preference probabilities, i.e,
$t(y|x)\propto\pi_{\texttt{ref}}(y|x)p(I=0|x,y)$
or in case of access to q-function or reward function
$t(y|x)\propto\pi_{\texttt{ref}}(y|x)\exp(\frac{-Q(x,y)}{\eta})$
Intuitively this objective gets closer to preferred distribution $q(y|x)$ and gets away from dis-preferred distribution $t(y|x)$ while staying close to reference policy.
In simplest case when we have access to rejected and/or accepted actions $\{{y_{a}}^{i}\sim\pi_{\operatorname*{\mathrm{ref}}},{y_{r}}^{i}\sim\pi_{\operatorname*{\mathrm{ref}}},x^{i}\sim\mu(x)\}_{i=1...N}$ (In case of Q-function we simply sample 2 actions and the action with higher Q-value is accepted and the one with lower q-value is rejected). The above general update rule will simplify to:
$\pi_{\operatorname*{\mathrm{imp}}}=\operatorname*{\mathrm{arg~{}max}}_{\theta}\alpha\Sigma_{i}\log\pi_{\theta}(y_{a}^{i}|x^{i})-(1-\alpha)\Sigma_{i}\log\pi_{\theta}(y_{r}^{i}|x^{i})-\beta\Sigma_{i}\operatorname*{\mathrm{KL}}(\pi_{\operatorname*{\mathrm{ref}}}|\pi_{\theta})$
where $\pi_{\theta}$ is initialized with $\pi_{\operatorname*{\mathrm{ref}}}$.
Appendix J Benchmarks
J.1 Control Suite
The DeepMind Control Suite (tunyasuvunakool2020) is a collection of benchmark tasks implemented in the MuJoCo simulator (TodorovET12mujoco). The suite includes a variety of embodiments of different complexity and action dimensionality. For each of these embodiments, there are multiple tasks implemented, each of which defines a single reward function. Example images for some of the domains are shown in Figure 2.
Figure 2: Example domains in the Control Suite. From left to right: Cartpole, Acrobot, Reacher, Manipulator, Cheetah, Humanoid
J.2 RGB Stacking
The RGB Stacking benchmark (lee2021pickandplace) is a robotics task involving a Rethink Sawyer robot arm outfitted with a Robotiq 2F-85 gripper, as well as a basket containing a number of parameterized geometric shapes in red, green and blue color. See Figure 3 for an illustration. The goal of this task is to have the robot arrange the shapes into varying arrangements, such as stacking one on top of another or building a tower. The policy only provides proprioception information and the images from three cameras surrounding the basket; there is no explicit tracking of the objects’ relative positions, or their identities. Specifically, the agent is provided the observations listed in Table 1. Thus the task’s challenge lies in forcing the agent to learn a control policy directly from vision, and recognizing which objects are in the workspace, since their different geometric properties demand different manipulation strategies.
Figure 3: Illustration of the simulated RGB Stacking domain. Top right corner: goal image (left) and current observation (right) for the front left camera, both as provided to the agent. Bottom: reward trace for the dense triple-stacking objective.
Figure 4: Comparison of PMPO and DPO for high-dimensional control tasks from the DeepMind Control Suite. We plot average reward over time of training (using 100 episodes for each evaluation). For each state we sample 4 responses from reference policy. We compare PMPO and DPO when both use only 2 samples for learning; the best and worst responses only (Accept[Best]&Reject[Worst]). We also compare PMPO and DPO when both use all 4 samples. For PMPO these would be top two as preferred generations and bottom two as dispreferred generations (Accept[Best]&Reject[Worst]). For DPO we create two sets of pairs, i.e., (best and worst) and (second best and second worst). Results show in general using two samples for learning and all 4 samples in learning yielded similar performance.
Modality Dimensions
Arm joint angles 7 x 3
Arm joint velocities 7 x 3
Arm joint torque 7 x 3
Gripper motor angle 1 x 3
Gripper motor velocity 1 x 3
Gripper grasp sensor 1 x 3
Cartesian tool center point pose 7 x 3
Cartesian wrist endpoint velocity 6 x 3
Cartesian wrist endpoint force 3 x 3
Cartesian wrist endpoint torque 3 x 3
Back left basket camera 80 x 80
Front left basket camera 80 x 80
Front right basket camera 80 x 80
Table 1: Observations given to the agent in the RGB Stacking benchmark. Note that for all observations except the camera images, a history of 3 time steps is provided, resulting in the x3 dimensionalities.
The task is implemented in the MuJoCo physics simulator (TodorovET12mujoco). Using the ground truth positions of the objects’ positions (which is not provided to the agent), several reward terms are calculated, including stacking two objects, stacking three objects, and building a pyramid, for each object permutation. Details of these rewards can be found in springenberg2024offlineactorcriticreinforcementlearning.
Appendix K Additional Experiments
In this section we evaluate the performance of PMPO and DPO in high-dimensional control tasks from the DeepMind Control Suite (Figure 4). For each state, we sample four responses from a reference policy. We compare PMPO and DPO under two conditions:
  1. 1.
    Two Samples: Both algorithms utilize only two samples for learning; the best and worst responses (Accept[Best] & Reject[Worst]).
  2. 2.
    Four Samples: Both algorithms utilize all four samples (Accept&Reject). For PMPO, the top two responses are used as preferred generations and the bottom two as dispreferred generations. For DPO, we create two sets of pairs for each prompt: (best and worst) and (second best and second worst).
Figure 4 indicates that using two samples for learning yields similar performance to using all four samples. Moreover, PMPO remains competitive with DPO when using pairs, while generally performing better when the policy is represented as Gaussian.